# Polymatroid Schubert varieties

### v1

@@ Abstract.
The lattice of flats $\mathcal{L}_{M}$ of a matroid $M$ is combinatorially well-behaved and, when $M$ is realizable, admits a geometric model in the form of a “matroid Schubert variety”. In contrast, the lattice of flats of a polymatroid exhibits many combinatorial pathologies and admits no similar geometric model.
We address this situation by defining the lattice $\mathcal{L}_{P}$ of “combinatorial flats” of a polymatroid $P$. Combinatorially, $\mathcal{L}_{P}$ exhibits good behavior analogous to that of $\mathcal{L}_{M}$: it is graded, determines $P$ when $P$ is simple, and is top-heavy. When $P$ is realizable over a field of characteristic 0, we show that $\mathcal{L}_{P}$ is modelled by a “polymatroid Schubert variety”.
Our work generalizes a number of results of Ardila-Boocher and Huh-Wang on matroid Schubert varieties; however, the geometry of polymatroid Schubert varieties is noticeably more complicated than that of matroid Schubert varieties. Many natural questions remain open.
@@ 1. Introduction
Let $M$ be a matroid of rank $d$ on an $N$-element set $E$, and let $\mathcal{L}_{M}$ be its poset of flats. Beloved combinatorial properties of $\mathcal{L}_{M}$ include:
@@
  • $\mathcal{L}_{M}$ is top-heavy: if $\mathcal{L}_{M}^{k}$ is the set of rank $k$ elements of $\mathcal{L}_{M}$, then $|\mathcal{L}_{M}^{k}|\leq|\mathcal{L}_{M}^{d-k}|$ when $2k\leq d$.
Top-heaviness was long conjectured [DW74, DW75], but proved only recently in [BHM+20b]. The approach of [BHM+20b] is inspired by [HW17], who proved top-heaviness when $M$ is the matroid associated to a linear subspace $V\subset\mathbb{K}^{N}$, with $\mathbb{K}$ a field. Central to [HW17] is the matroid Schubert variety $Y_{V}$, defined in [AB16] as the closure of $V\subset\mathbb{K}^{N}$ in $(\mathbb{P}^{1})^{n}=(\mathbb{K}\cup\infty)^{N}$.
The matroid Schubert variety $Y_{V}$ may be thought of as an algebro-geometric counterpart of $\mathcal{L}_{M}$, fully capturing the combinatorics of $M$. To see why, first observe that the additive action of $\mathbb{K}$ extends to an action on $\mathbb{P}^{1}$, fixing $\infty$. Hence, $\mathbb{K}^{n}$ acts on $(\mathbb{P}^{1})^{n}$. The orbits of this action are of the form
@@
  • the class of $\overline{Y_{V}\cap O_{S}}$ in the Chow ring of $(\mathbb{P}^{1})^{n}$ is of the form
@@
We propose to remedy these pathologies with a new notion of flats for polymatroids.
@@
Let $P$ be a polymatroid. There exists a poset of combinatorial flats $\mathcal{L}_{P}$ such that
@@
    there is a unique simple polymatroid $P^{\mathrm{sim}}$ such that $\mathcal{L}_{P}\cong\mathcal{L}_{P^{\mathrm{sim}}}$, and
@@
    $\mathcal{L}_{P}$ is top-heavy.
When $P$ is realized by a subspace arrangement $V_{1},\ldots,V_{N}$, the lattice $\mathcal{L}_{P}$ is the poset of flats of a collection of general flags with minimal elements $V_{1},\ldots,V_{N}$ (see 1.2). For arbitrary polymatroids, $\mathcal{L}_{P}$ may be constructed as follows [^1] The definitions of “rank” and “combinatorial flat” given here are different from, but equivalent to, the definitions given in Section 3.1. The equivalence is explained in 3.6.. Let $\operatorname{rk}_{P}$ be the rank function of $P$, and let $\mathbf{e}_{1},\ldots,\mathbf{e}_{N}$ be the standard basis of ${\mathbb{Z}}^{N}$. Choose $\mathbf{n}=(n_{1},\ldots,n_{N})\in\mathbb{Z}_{\geq 0}^{N}$, such that $n_{i}\geq\operatorname{rk}_{P}(i)$. Recall that an independent multiset of $P$ is $\mathbf{b}=(b_{1},\ldots,b_{N})\in\mathbb{Z}_{\geq 0}^{N}$ such that $\sum_{i\in A}b_{i}\leq\operatorname{rk}_{P}(A)$ for all $A\subset E$. A basis for a multiset $\mathbf{s}=(s_{1},\ldots,s_{N})\in\mathbb{Z}_{\geq 0}^{N}$ is an independent multiset $\mathbf{b}=(b_{1},\ldots,b_{N})$, maximal among those contained in $\mathbf{s}$. Define the rank of $\mathbf{s}$ by $\operatorname{rk}_{P}(\mathbf{s}):=\sum_{i\in E}b_{i}$ for any basis $\mathbf{b}$2223.5 explains well-definedness of rank. We write $\operatorname{rk}_{P}(\cdot)$ for ranks of both sets and multisets; this will never cause confusion., and call $\mathbf{s}$ a combinatorial flat [^3] This definition resembles the related notion of “flats for matricubes” [AG24]. See 3.8 for comparison of these notions. if $\operatorname{rk}_{P}(\mathbf{s}+\mathbf{e}_{i})>\operatorname{rk}_{P}(\mathbf{e}_{i})$ for all $i$ such that $s_{i}<n_{i}$. When $P$ is a matroid, $\mathcal{L}_{P}$ is its usual lattice of flats.
Let $\mathcal{L}_{P}$ be the set of combinatorial flats, ordered by inclusion. We will see in Section 3.4 that the poset $\mathcal{L}_{P}$ does not depend on $\mathbf{n}$ (although the underlying set of combinatorial flats does; we write $\mathcal{L}_{P,\mathbf{n}}$ when we wish to emphasize this).
@@
Let $\mathbf{n}=(1,1,1,2)$. The lattice $\mathcal{L}_{P}\cong\mathcal{L}_{P,\mathbf{n}}$ of combinatorial flats is both graded and top-heavy. Elements of $\mathcal{L}_{P}$ correspond to flats of the arrangement $\{V_{1},V_{2},V_{3},V_{4},V_{4,1}\}$, where $V_{4,1}$ is a general plane containing $V_{4}$. There is a lattice homomorphism embedding the lattice of flats of $P$ into $\mathcal{L}_{P}$ (Fig. 2). This construction is explained in full generality in 3.7.
@@
Unlike the usual poset of flats of $P$, which has no known algebro-geometric realization, $\mathcal{L}_{P}$ can be realized as the poset of affine cells of a projective variety whenever $P$ is realizable.
Explicitly, work of Hassett-Tschinkel [HT99] (reviewed in Section 4.1) shows that the additive action of $\mathbb{K}^{\mathbf{n}}:=\mathbb{K}^{n_{1}}\times\cdots\times\mathbb{K}^{n_{N}}$ on itself may be extended to an action on $\mathbb{P}^{\mathbf{n}}:=\mathbb{P}^{n_{1}}\times\cdots\times\mathbb{P}^{n_{N}}$. The action partitions $\mathbb{P}^{\mathbf{n}}$ into finitely many orbits $O_{\mathbf{s}}$, corresponding to multisets $\mathbf{s}\leq\mathbf{n}$. Under the assumption that $V$ is “polymatroid general”, a mild genericity condition described in Section 4.6 that constrains the intersection of $V$ with certain coordinate subspaces, we prove the following theorem.
@@
If $\mathbb{K}=\mathbb{C}$, we also give a presentation for the singular cohomology ring of $Y_{V}$, generalizing a result of [HW17]. By 1.1i, any two combinatorial flats $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{L}_{P}$ have a least upper bound $\mathbf{s}\vee\mathbf{s}^{\prime}\in\mathcal{L}_{P}$.
@@
where $\deg y_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. There are constants $c_{\mathbf{b}}$ for each basis $\mathbf{b}$ such that multiplication is given by the formula
$c_{\mathbf{b}}c_{\mathbf{b}^{\prime}}y_{\mathbf{s}}y_{\mathbf{s}^{\prime}}=\begin{cases}c_{\mathbf{b}+\mathbf{b}^{\prime}}y_{\mathbf{s}\vee\mathbf{s}^{\prime}},&\text{if $\operatorname{rk}_{P}(\mathbf{s}\vee\mathbf{s}^{\prime})=\operatorname{rk}_{P}(\mathbf{s})+\operatorname{rk}_{P}(\mathbf{s}^{\prime})$}\\ 0,&\text{otherwise,}\end{cases}$
where $\mathbf{b}$ and $\mathbf{b}^{\prime}$ are bases for $\mathbf{s}$ and $\mathbf{s}^{\prime}$, respectively, such that $\mathbf{b}+\mathbf{b}^{\prime}$ is a basis for $\mathbf{s}\vee\mathbf{s}^{\prime}$.
1.4. Acknowledgements
We thank Lucas Gierczak for helping to sort out the relationship between the “matricubes” of [AG24] and the objects of the present work (see 3.8). We are all partially supported by the Simons Foundation; the first author is also partially supported by NSF grant DMS-2039316. We are grateful to the Institute for Advanced Study (IAS). The second and third authors also thank the Beijing International Center for Mathematical Research (BICMR) for their hospitality and support during the preparation of this paper.
@@ 1.5. Organization
We review polymatroids, multisymmetric matroids, and operations on them in Section 2, before comencing study of combinatorial flats in Section 3.
@@
Section 4 concerns polymatroid Schubert varieties. Foundational work of Hassett and Tschinkel [HT99] is reviewed in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section 4.4, then compute the support of its Chow class in Section 4.5. In Section 4.6, we discuss “polymatroid generality”, a condition on $V$ that guarantees $Y_{V}$ consists of finitely many $V$-orbits. Assuming this condition, we identify strata of $Y_{V}$ in Section 4.7, culminating with a proof of 1.3. In Section 4.8, we prove 1.4.
At the last, we pose a large number of questions about polymatroid Schubert varieties in Section 5. Both algebraic and geometric aspects of polymatroid Schubert varieties seem markedly more complicated than those of matroid Schubert varieties, yet examples suggest a good theory lies in wait. Interested parties are encouraged to seek their fortune.
@@ 2. Polymatroids
Let $\widetilde{E}=E_{1}\sqcup\cdots\sqcup E_{N}$ be a finite set. A multisymmetric matroid is a matroid $M$ on $\widetilde{E}$ whose rank function commutes with the natural action of $\mathfrak{S}:=\prod_{i=1}^{N}\mathfrak{S}_{E_{i}}$. If $F$ is a flat of $M$ and $\sigma\in\mathfrak{S}$, then $\sigma\cdot F$ is also a flat of $M$. The geometric part of a subset $A\subset\widetilde{E}$ is $A^{\mathrm{geo}}:=\cap_{\sigma\in\mathfrak{S}}\sigma(A)$, and $A$ is geometric if $A=A^{\mathrm{geo}}$. If $F$ is a flat, then $F^{\mathrm{geo}}$ is also a flat. Two key properties of multisymmetric matroids follow.
@@
Multisymmetric matroids and polymatroids are related by a much-rediscovered [Hel72, Ngu86, Lov77, McD75, BCF23, CHL+20] recipe. A cage for a polymatroid $P$ on $E$ is $\mathbf{n}\in\mathbb{Z}_{\geq 0}^{E}$ such that $\operatorname{rk}(i)\leq n_{i}$ for all $i\in E$. The pair $(P,\mathbf{n})$ is a caged polymatroid [^4] This terminology originates in [EL23]..
@@
Equivalently, $B\subset\widetilde{E}$ is a basis of $\widetilde{P}$ if and only if $\big{(}|B\cap E_{i}|\big{)}_{i=1}^{N}$ is a basis of $P$.
@@
$\geq|\pi_{\overline{A}}(\mathbf{b})|\geq|\pi_{A}(\mathbf{b})|,$ so $\mathbf{b}$ is a basis of $\widetilde{P}$.
@@
It suffices to give a proof for $F:=\overline{S}$. Since $\cup_{i\in F}E_{i}$ is a geometric set of $\widetilde{P}$, the truncation $T_{\cup_{i\in F}E_{i}}\widetilde{P}$ is also a multisymmetric matroid on $\widetilde{E}$. For $A\subset E$,
$\operatorname{rk}_{T_{\cup_{i\in F}E_{i}}\widetilde{P}}(\cup_{i\in A}E_{i})$ =$=\begin{cases}\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})-1,&\text{if $\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})=\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A\cup F}E_{i})$}\\ \operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i}),&\text{otherwise}\end{cases}$
=$=\begin{cases}\operatorname{rk}_{P}(A)-1,&\text{if $\operatorname{rk}_{P}(A)=\operatorname{rk}_{P}(A\cup F)$}\\ \operatorname{rk}_{P}(A),&\text{otherwise}\end{cases}$
$=\operatorname{rk}_{T_{F}P}(A).$
Consequently, $\widetilde{T_{F}P}=T_{\cup_{i\in F}E_{i}}\widetilde{P}$ by 2.5. ∎
@@
$\operatorname{rk}_{\widetilde{P}\setminus j}((\cup_{k\in A}E_{k})\setminus j)=\operatorname{rk}_{\widetilde{P}}((\cup_{k\in A}E_{k})\setminus j)=\operatorname{rk}_{\widetilde{P}}((\cup_{k\in A}E_{k}))=\operatorname{rk}_{P}(A)=\operatorname{rk}_{R_{i}P}(A).$

### v2

# Combinatorial flats and Schubert varieties of subspace arrangements

@@ Abstract.
The lattice of flats $\mathcal{L}_{M}$ of a matroid $M$ is combinatorially well-behaved and, when $M$ is realizable, admits a geometric model in the form of a “Schubert variety of hyperplane arrangement”. In contrast, the lattice of flats of a polymatroid exhibits many combinatorial pathologies and admits no similar geometric model.
We address this situation by defining the lattice $\mathcal{L}_{P}$ of “combinatorial flats” of a polymatroid $P$. Combinatorially, $\mathcal{L}_{P}$ exhibits good behavior analogous to that of $\mathcal{L}_{M}$: it is graded, determines $P$ when $P$ is simple, and is top-heavy. When $P$ is realizable over a field of characteristic 0, we show that $\mathcal{L}_{P}$ is modeled by “the Schubert variety of a subspace arrangement”.
Our work generalizes a number of results of Ardila-Boocher and Huh-Wang on Schubert varieties of hyperplane arrangements; however, the geometry of Schubert varieties of subspace arrangements is noticeably more complicated than that of Schubert varieties of hyperplane arrangements. Many natural questions remain open.
@@ 1. Introduction
Let $M$ be a matroid of rank $d$ on an $N$-element set $E$, and let $\mathcal{L}_{M}$ be its poset of flats. Write $\mathcal{L}_{M}^{k}$ for the set of flats of rank $k$. Beloved combinatorial properties of $\mathcal{L}_{M}$ include:
@@
  • $|\mathcal{L}_{M}^{k}|\leq|\mathcal{L}_{M}^{k+1}|$ when $k\leq d/2n$.
  • $\mathcal{L}_{M}$ is top-heavy: $|\mathcal{L}_{M}^{k}|\leq|\mathcal{L}_{M}^{d-k}|$ when $k\leq d/2$.
Top-heaviness was long conjectured [DW74, DW75], but proved only recently in [BHM+20]. The approach of [BHM+20] is inspired by [HW17], who proved top-heaviness when $M$ is the matroid associated to a linear subspace $V\subset\mathbb{K}^{N}$, with $\mathbb{K}$ a field. Central to [HW17] is the Schubert variety of hyperplane arrangement $Y_{V}$, defined in [AB16] as the closure of $V\subset\mathbb{K}^{N}$ in $(\mathbb{P}^{1})^{N}=(\mathbb{K}\cup\infty)^{N}$.
The Schubert variety of hyperplane arrangement $Y_{V}$ may be thought of as an algebro-geometric counterpart of $\mathcal{L}_{M}$, fully capturing the combinatorics of $M$. To see why, first observe that the additive action of $\mathbb{K}$ on itself extends to an action on $\mathbb{P}^{1}$, fixing $\infty$. Hence, $\mathbb{K}^{N}$ acts on $(\mathbb{P}^{1})^{N}$. The orbits of this action are of the form
@@
  • the class of $\overline{Y_{V}\cap O_{S}}$ in the Chow ring of $(\mathbb{P}^{1})^{n}$ is
@@
To remedy these pathologies, we introduce the lattice of combinatorial flats $\mathcal{L}_{P}$ of a polymatroid. When $P$ is a matroid, $\mathcal{L}_{P}$ is the usual lattice of flats. When $P$ is realized by a subspace arrangement $V_{1},\ldots,V_{N}$, the lattice $\mathcal{L}_{P}$ is the poset of flats of a collection of general flags with minimal elements $V_{1},\ldots,V_{N}$ (see 1.2).
For arbitrary polymatroids, $\mathcal{L}_{P}$ may be constructed as follows [^1] The definitions of “rank” and “combinatorial flat” given here are different from, but equivalent to, the definitions given in Section 3.1. The equivalence is explained in 3.6.. Let $\operatorname{rk}_{P}$ be the rank function of $P$, and let $\mathbf{e}_{1},\ldots,\mathbf{e}_{N}$ be the standard basis of ${\mathbb{Z}}^{N}$. Choose $\mathbf{n}=(n_{1},\ldots,n_{N})\in\mathbb{Z}_{\geq 0}^{N}$, such that $n_{i}\geq\operatorname{rk}_{P}(i)$. Recall that an independent multiset of $P$ is $\mathbf{b}=(b_{1},\ldots,b_{N})\in\mathbb{Z}_{\geq 0}^{N}$ such that $\sum_{i\in A}b_{i}\leq\operatorname{rk}_{P}(A)$ for all $A\subset E$. A basis for a multiset $\mathbf{s}=(s_{1},\ldots,s_{N})\in\mathbb{Z}_{\geq 0}^{N}$ is an independent multiset $\mathbf{b}=(b_{1},\ldots,b_{N})$, maximal among those contained in $\mathbf{s}$. Define the rank of $\mathbf{s}$ by $\operatorname{rk}_{P}(\mathbf{s}):=\sum_{i\in E}b_{i}$ for any basis $\mathbf{b}$2223.5 explains well-definedness of rank. We write $\operatorname{rk}_{P}(\cdot)$ for ranks of both sets and multisets; this will never cause confusion., and call $\mathbf{s}$ a combinatorial flat [^3] This definition resembles the related notion of “flats for matricubes” [AG24]. See 3.8 for comparison of these notions. if $\operatorname{rk}_{P}(\mathbf{s}+\mathbf{e}_{i})>\operatorname{rk}_{P}(\mathbf{s})$ for all $i$ such that $s_{i}<n_{i}$.
We define $\mathcal{L}_{P}$ to be the set of combinatorial flats, ordered by inclusion. We will see in Section 3.4 that $\mathcal{L}_{P}$ does not depend on $\mathbf{n}$ (although the underlying set of combinatorial flats does; we write $\mathcal{L}_{P,\mathbf{n}}$ when we wish to emphasize this).
@@
Let $P$ be a polymatroid. The poset of combinatorial flats satisfies:
@@
    there is a unique simple polymatroid $P^{\mathrm{sim}}$ such that $\mathcal{L}_{P}\cong\mathcal{L}_{P^{\mathrm{sim}}}$,
@@
    $|\mathcal{L}_{P}^{k}|\leq|\mathcal{L}_{P}^{k+1}|$ when $k\leq d/2$, and
  4. (iv)
    $\mathcal{L}_{P}$ is top-heavy: $|\mathcal{L}_{P}^{k}|\leq|\mathcal{L}_{P}^{d-k}|$ when $k\leq d/2$.
@@
On the other hand, $\mathcal{L}_{P}\cong\mathcal{L}_{P,\mathbf{n}}$ (with $\mathbf{n}=(1,1,1,2)$) is both graded and top-heavy (Fig. 2). Elements of $\mathcal{L}_{P}$ correspond to flats of the arrangement $\{V_{1},V_{2},V_{3},V_{4},V_{4,1}\}$, where $V_{4,1}$ is a general plane containing $V_{4}$. There is a lattice homomorphism embedding the lattice of flats of $P$ into $\mathcal{L}_{P}$. This construction is explained in full generality in 3.7.
@@
Lattices of combinatorial flats can be axiomatized in a manner that closely resembles the well-known definition of geometric lattices (3.20). Consequently, they constitute a new cryptomorphic definition of simple polymatroids.
@@
Unlike the usual poset of flats of $P$, $\mathcal{L}_{P}$ can be realized as the poset of affine cells of a projective variety whenever $P$ is realizable over a field of characteristic zero.
Work of Hassett-Tschinkel [HT99] (reviewed in Section 4.1) shows that there exists action of the additive group $\mathbb{K}^{\mathbf{n}}:=\mathbb{K}^{n_{1}}\times\cdots\times\mathbb{K}^{n_{N}}$ on $\mathbb{P}^{\mathbf{n}}:=\mathbb{P}^{n_{1}}\times\cdots\times\mathbb{P}^{n_{N}}$, such that $\mathbb{P}^{\mathbf{n}}$ is partitioned into finitely many orbits $O_{\mathbf{s}}$, corresponding to multisets $\mathbf{s}\leq\mathbf{n}$. Under the assumption that $V$ is “polymatroid general”, a mild genericity condition described in Section 4.6 that constrains the intersection of $V$ with certain coordinate subspaces, we define $Y_{V}$ to be the $V$-orbit closure of the origin in $\mathbb{P}^{\mathbf{n}}$ under the Hasset-Tschinkel action of $\mathbb{K}^{\mathbf{n}}$ on $\mathbb{P}^{\mathbf{n}}.$
@@
If $\mathbb{K}=\mathbb{C}$, we also give a presentation for the singular cohomology ring of $Y_{V}$, generalizing a result of [HW17].
Definition 1.4.
Let $P$ be a polymatroid with ground set $E=\{1,\ldots,N\}$. Given two combinatorial flats $\mathbf{s}$ and $\mathbf{s^{\prime}}$ spanned by independent multisets $\mathbf{b}$ and $\mathbf{b^{\prime}}$ respectively, we define their cup product as
$(\mathbf{s}\smile\mathbf{s^{\prime}}):=\overline{\left(\min\{\mathbf{b}(i)+\mathbf{b^{\prime}}(i),\operatorname{rk}_{P}(i)\}\right)_{i=1}^{N}}.$
Remark 1.5.
Equivalently, in terms of Definition 3.1, $\mathbf{s}\smile\mathbf{s^{\prime}}$ is the combinatorial flat represented by any maximal element of $\{F\vee F^{\prime}\}$ where $F$ and $F^{\prime}$ range over flats of the multisymmetric lift $\widetilde{P}$ which represent $\mathbf{s}$ and $\mathbf{s^{\prime}}$ respectively. Notice that the product of symmetric groups $\mathfrak{S}:=\prod_{i=1}^{N}\mathfrak{S}_{E_{i}}$ acts transitively on the maximal elements of $\{F\vee F^{\prime}\}$. Therefore $\mathbf{s}\smile\mathbf{s^{\prime}}$ is well defined, independent of the choices of $\mathbf{b}$ and $\mathbf{b^{\prime}}$.
Example 1.6.
Suppose that $N=1$, and $\operatorname{rk}_{P}(\{1\})=n$. Then the lattice $\mathcal{L}_{P}$ of combinatorial flats is isomorphic to the chain of length $n$. Any combinatorial flat $i\in[n]$ is also independent in $P$, so the cup product of combinatorial flats is given as
$i\smile j:=\overline{\min(i+j,n)}=\min(i+j,n).$
$\diamondsuit$
@@
where $\deg y_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. There are positive constants $c_{\mathbf{b}}$ associated to bases $\mathbf{b}$ of combinatorial flats such that multiplication is given by the formula
$c_{\mathbf{b}}c_{\mathbf{b}^{\prime}}y_{\mathbf{s}}y_{\mathbf{s}^{\prime}}=\begin{cases}c_{\mathbf{b}+\mathbf{b}^{\prime}}y_{\mathbf{s}\smile\mathbf{s}^{\prime}},&\text{if $\operatorname{rk}_{P}(\mathbf{s}\smile\mathbf{s}^{\prime})=\operatorname{rk}_{P}(\mathbf{s})+\operatorname{rk}_{P}(\mathbf{s}^{\prime})$}\\ 0,&\text{otherwise,}\end{cases}$
where $\mathbf{b}$ and $\mathbf{b}^{\prime}$ are bases for $\mathbf{s}$ and $\mathbf{s}^{\prime}$, respectively, such that $\mathbf{b}+\mathbf{b}^{\prime}$ is a basis for $\mathbf{s}\smile\mathbf{s}^{\prime}$.
@@ 1.4. Organization
We review polymatroids, multisymmetric matroids, and operations on them in Section 2, before commencing study of combinatorial flats in Section 3.
@@
Section 4 concerns Schubert varieties of subspace arrangements. Foundational work of Hassett and Tschinkel [HT99] is reviewed in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section 4.4, then compute the support of its Chow class in Section 4.5. In Section 4.6, we discuss “polymatroid generality”, a condition on $V$ that guarantees $Y_{V}$ consists of finitely many $V$-orbits. Assuming this condition, we identify strata of $Y_{V}$ in Section 4.7, culminating with a proof of 1.3. In Section 4.8, we prove 1.7.
At the last, we pose a large number of questions about Schubert varieties of subspace arrangements in Section 5. Both algebraic and geometric aspects of Schubert varieties of subspace arrangements seem markedly more complicated than those of Schubert varieties of hyperplane arrangements, yet examples suggest a good theory lies in wait. Interested parties are encouraged to seek their fortune.
@@ 1.5. Notation
1.6. Acknowledgements
We thank Lucas Gierczak for helping to sort out the relationship between the “matricubes” of [AG24] and the objects of the present work (see 3.8), and Matt Larson for helpful comments on our draft. We are all partially supported by the Simons Foundation. The first author is partially supported by NSF grant DMS-2039316, and the last author is partially supported by NSF grant DMS-1926686. We are grateful to the Institute for Advanced Study (IAS). The second and third authors also thank the Beijing International Center for Mathematical Research (BICMR) for their hospitality and support during the preparation of this paper.
@@ 2. Polymatroids
Let $\widetilde{E}=E_{1}\sqcup\cdots\sqcup E_{N}$ be a finite set. A multisymmetric matroid is a matroid $M$ on $\widetilde{E}$ whose rank function is invariant under the natural action of $\mathfrak{S}:=\prod_{i=1}^{N}\mathfrak{S}_{E_{i}}$. If $F$ is a flat of $M$ and $\sigma\in\mathfrak{S}$, then $\sigma\cdot F$ is also a flat of $M$. The geometric part of a subset $A\subset\widetilde{E}$ is $A^{\mathrm{geo}}:=\cap_{\sigma\in\mathfrak{S}}\sigma(A)$, and $A$ is geometric if $A=A^{\mathrm{geo}}$. If $F$ is a flat, then $F^{\mathrm{geo}}$ is also a flat. Two key properties of multisymmetric matroids follow.
@@
Multisymmetric matroids and polymatroids are related by a much-rediscovered [Hel72, Ngu86, Lov77, McD75, BCF23, CHL+20] recipe. A cage for a polymatroid $P$ on $E$ is $\mathbf{n}\in\mathbb{Z}_{\geq 0}^{E}$ such that $\operatorname{rk}(i)\leq n_{i}$ for all $i\in E$. The pair $(P,\mathbf{n})$ is a caged polymatroid [^4] This terminology originates in [EL24]..
@@
Equivalently, $B\subset\widetilde{E}$ is a basis of $\widetilde{P}$ if and only if $\big{(}|B\cap E_{1}|,\ldots,|B\cap E_{N}|\big{)}$ is a basis of $P$.
@@
$\geq|\pi_{\overline{A}}(\mathbf{b})|\geq|\pi_{A}(\mathbf{b})|,$ so $\mathbf{b}$ is a basis of $\widetilde{P}$. $\qquad\qed$
@@
Example 2.11.
Suppose that $\mathbf{n}=(3)$ and that $P$ is the polymatroid of rank 2 on one element, realized by a line $V\subset\mathbb{K}^{\mathbf{n}}$. Every line in $\mathbb{K}^{\mathbf{n}}$ realizes $P$ because every line is codimension 2. However, different lines can have different matroids, e.g. a general line realizes a uniform matroid of rank 1 on 3 elements, while each of the three coordinate lines realizes a boolean matroid on 1 element with two loops adjoined. $\diamondsuit$
@@
Since $\cup_{i\in S}E_{i}$ is a geometric set of $\widetilde{P}$, the truncation $T_{\cup_{i\in S}E_{i}}\widetilde{P}$ is also a multisymmetric matroid on $\widetilde{E}$. For $A\subset E$,
$\operatorname{rk}_{T_{\cup_{i\in S}E_{i}}\widetilde{P}}(\cup_{i\in A}E_{i})$ =$=\begin{cases}\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})-1,&\text{if $\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i})=\operatorname{rk}_{\widetilde{P}}(\cup_{i\in A\cup S}E_{i})$}\\ \operatorname{rk}_{\widetilde{P}}(\cup_{i\in A}E_{i}),&\text{otherwise}\end{cases}$
=$=\begin{cases}\operatorname{rk}_{P}(A)-1,&\text{if $\operatorname{rk}_{P}(A)=\operatorname{rk}_{P}(A\cup S)$}\\ \operatorname{rk}_{P}(A),&\text{otherwise}\end{cases}$
$=\operatorname{rk}_{T_{S}P}(A).$
Consequently, $\widetilde{T_{S}P}=T_{\cup_{i\in S}E_{i}}\widetilde{P}$ by 2.5. ∎
@@
$\operatorname{rk}_{\widetilde{P}\setminus j}((\cup_{k\in A}E_{k})\setminus j)=\operatorname{rk}_{\widetilde{P}}((\cup_{k\in A}E_{k})\setminus j)=\operatorname{rk}_{\widetilde{P}}((\cup_{k\in A}E_{k}))=\operatorname{rk}_{P}(A)=\operatorname{rk}_{R_{i}P}(A).\qed$
We defer discussion of realizing reductions to 3.17, where we will see it in action en route to proving 1.1(ii).